Cannot use calculator in class what is the best way to work out (0.08)4power
Answer by stanbon(75887)

Cannot use calculator in class what is the best way to work out (0.08)4power
--------
0.08^4 = (8/100)^4 = (2/25)^4 = (2*2*2*2)/(25*25*25*25) = 16/390625
=======
Cheers,
Stan H.